# Descrip3D: Enhancing Large Language Model-based 3D Scene Understanding with Object-Level Text Descriptions

### v1

@@ 4.2 Implementation Details
Table 1: Performance on ScanQA [2], SQA3D [34], and ScanRefer [4]. “Expert models” are tailored for specific tasks using task-oriented heads, while “LLM-based models” are designed for general instructions and responses. Descrip3D achieves the highest performance across all benchmarks, outperforming prior LLM-based methods by leveraging relational textual descriptions through a dual-level integration for more precise and context-aware reasoning.
4.3 Performance Comparison
Scene-level Question Answering.
We evaluate our model on ScanQA [2] and SQA3D [34], both of which require comprehensive scene-level understanding and precise object-grounded reasoning. As shown in Table 1, our method achieves the highest scores across all major metrics compared with previous expert models and LLM-based methods, highlighting the strength of our dual-level integrated relational descriptions in enhancing context understanding. Notably, CIDEr, which emphasizes content relevance, shows a significant gain of 5.9 over PQ3D and 6.0 over Chat-Scene, confirming that our model generates more informative and relevant answers. On the SQA3D dataset, our method attains the highest EM and EM-R, again outperforming existing models. The strong results across both benchmarks demonstrate the effectiveness of enriching 3D object representations with object-level textual descriptions. As a general-purpose model, our model consistently outperforms task-specific systems, demonstrating strong versatility without the need for task-dependent architectures. Compared to knowledge-graph-based approaches such as 3DGraphLLM, our model achieves better alignment with language and context, suggesting that lightweight textual descriptions offer a more direct and interpretable semantic grounding. Overall, the results confirm that injecting fine-grained natural language descriptions of objects into the scene representation, through dual-level integration, significantly enhances the LLM’s ability to handle 3D visual question answering tasks.
3D Visual Grounding.
We evaluate 3D visual grounding on the ScanRefer [4] dataset. As shown in Table 1, our method achieves the best performance among all methods. Compared to strong expert models like ConcreteNet and 3D-VisTA, our model surpasses them by 5.3 and 6.3 at the stricter 0.5 IoU threshold, respectively. These expert models are trained with tailored 3D architectures and task-specific objectives, highlighting the strength of our unified and generalizable approach. Among LLM-based methods, our model also clearly outperforms Chat-Scene and 3DGraphLLM. The improvement over Chat-Scene, which uses only 2D and 3D embeddings, validates the great benefit of incorporating relational object-level textual descriptions through a dual-level integration approach.
Multi-Object 3D Visual Grounding.
We evaluate our method on Multi3DRefer [58], a challenging benchmark requiring models to resolve complex multi-object visual grounding within 3D scenes. Unlike traditional single-object grounding, this task demands fine-grained relational reasoning among multiple entities. As shown in Table 2, our method achieves the best performance, outperforming both expert models and LLM-based baselines such as Chat-Scene and Grounded 3D-LLM.
@@ 4.3 Performance Comparison
3DVG-Transformer [59] - 25.5
3DJCG [3] - 26.6
D3Net [5] - 32.2
M3DRef-CLIP [58] 42.8 38.4
PQ3D [63] - 50.1
@@
Grounded 3D-LLM [10] 44.7 40.8
Chat-Scene [22] 57.1 52.4
@@
Table 2: Performance on Multi3DRefer [58]. Descrip3D achieves the best performance, with notable gains in multi-object reasoning due to its dual-level relational text modeling.
3D Captioning.
We evaluate our method on the Scan2Cap [11] benchmark, which focuses on generating natural language descriptions for 3D objects in complex indoor scenes. As shown in Table 3, our method achieves the highest CIDEr score, indicating that the generated captions are more informative and semantically aligned with the ground truth. While our BLEU-4 score is lower than some other models, this metric emphasizes exact n-gram overlap and can penalize variation in phrasing, particularly in models like ours that incorporate diverse relational language. Importantly, Descrip3D is trained as a unified model across multiple tasks without task-specific tuning, unlike expert models tailored for captioning. Despite this, it still delivers a strong overall performance, suggesting that our relational text descriptions effectively enhance semantic grounding and generalization across diverse scene-language tasks.
@@
Scan2Cap [11] 35.2 22.4
3DJCG [3] 49.5 31.0
3D-VLP [27] 54.9 32.3
3D-VisTA [62] 66.9 34.0
Vote2Cap-DETR++ [9] 67.6 37.1
@@
LL3DA [8] 65.2 36.8
LEO [23] 68.4 36.9
Grounded 3D-LLM [10] 70.2 35.0
Chat-Scene [22] 77.1 36.3
@@
Table 3: Performance on Scan2Cap [11]. Descrip3D achieves the best C@0.5 performance by leveraging richer object-level relational context.
Qualitative results.
Figure 3 presents qualitative comparisons between our method and Chat-Scene on two representative tasks: 3D question answering (Figure 3(a)) and 3D visual grounding (Figure 3(b)). For question answering, our model demonstrates a stronger ability to understand relational context and resolve spatial references. In comparison, Chat-Scene demonstrably struggles more with object relationships. For visual grounding, our model more accurately identifies target objects based on dual-level integrated complex natural language descriptions that involve appearance and relative position. In contrast, Chat-Scene frequently fails to disambiguate between visually or semantically similar candidates. These results highlight the benefit of injecting detailed object-level descriptions into the language model through a dual-level approach, enabling more precise and interpretable multi-modal reasoning.
(a) Qualitative Comparison of Question Answering.
(b) Qualitative Comparison of Grounding.
In summary, our method achieves strong and consistent results across the five benchmarks, demonstrating its effectiveness in different tasks.
4.4 Ablation Studies
To evaluate the impact of our proposed object-level text descriptions and the dual-level integration strategies, we conduct a series of ablation experiments across key components of our model.
Effect of Object Reference Style in Prompt-Level Description.
We examine how the choice of object reference expression in prompt-level text injection affects performance. As shown in Table 4, using raw object names can lead to suboptimal performance due to ambiguity when multiple instances of the same category are present. Appending object IDs to names may confuse the model further, failing to improve performance and potentially introducing redundancy. In contrast, using object IDs alone yields the best results across all benchmarks. This strategy eliminates ambiguity and ensures precise alignment between the referenced objects in the query and their corresponding descriptions.
@@ 4.4 Ablation Studies
Effect of Object Descriptions and Fusion Location.
We study the impact of incorporating object-level textual descriptions at different points in the model. As shown in Table 5, injecting textual descriptions solely as embeddings improves grounding and question answering by enriching object semantics. Injecting them into the prompt also boosts performance on tasks like ScanQA and Scan2Cap. Combining both yields the best results on most datasets, highlighting their complementarity. However, on Multi3DRefer, this strategy slightly underperforms, possibly due to redundant signals in complex scenes. Overall, the dual-level integration enhances both spatial grounding and semantic reasoning.
@@
5 Conclusion and Future Work

### v2

@@ 4 Experiments
Table 1: Performance on ScanQA [azuma2022scanqa], SQA3D [ma2022sqa3d], and ScanRefer [chen2020scanrefer]. “Expert models” are tailored for specific tasks using task-oriented heads, while “LLM-based models” are designed for general instructions and responses. Descrip3D achieves the highest performance across all benchmarks, outperforming prior LLM-based methods by leveraging relational textual descriptions through a dual-level integration for more precise and context-aware reasoning.
(a) Qualitative Comparison of Question Answering.
(b) Qualitative Comparison of Grounding.
4.1 Datasets and Metrics
Datasets.
To evaluate generalizability, we conduct experiments on five benchmarks spanning four major 3D vision-language tasks: ScanRefer [chen2020scanrefer] for single-object grounding, Multi3DRefer [zhang2023multi3drefer] for compositional multi-object grounding, Scan2Cap [chen2021scan2cap] for 3D-aware captioning, and ScanQA [azuma2022scanqa] plus SQA3D [ma2022sqa3d] for scene-level QA. All are derived from ScanNet [dai2017scannet], a richly annotated dataset of 1,513 indoor scenes with point clouds, RGB images, and camera poses. This shared backbone provides consistent semantics across tasks and enables multi-task evaluation. In our experiments, we strictly follow the official splits provided by each benchmark (ScanRefer, Multi3DRefer, Scan2Cap, ScanQA, and SQA3D). All are built on ScanNet, where splits are defined at the scene level, ensuring disjoint train/val/test sets. Thus, evaluation scenes and objects are never partially seen during training, eliminating data leakage.
Metrics.
We use standard evaluation protocols from prior work [huang2024chat]. For ScanRefer [chen2020scanrefer], we report thresholded accuracies Acc@0.25 and Acc@0.5, which assess whether the predicted object bounding box has an IoU with the ground truth exceeding 0.25 or 0.5. For Multi3DRefer [zhang2023multi3drefer], which involves grounding multiple targets, we use the F1 score at IoU thresholds of 0.25 and 0.5. For the captioning task Scan2Cap [chen2021scan2cap], we adopt CIDEr@0.5 and BLEU-4@0.5, integrating captioning quality with spatial alignment via IoU. For visual question answering, ScanQA [azuma2022scanqa] is evaluated using CIDEr and BLEU-4, while SQA3D [ma2022sqa3d] is evaluated using Exact Match (EM) and its refined variant EM-R as proposed in LEO [huang2023embodied].
3D 2D Text ScanRefer Multi3DRef Scan2Cap ScanQA SQA3D
Acc@0.5 F1@0.5 C@0.5 CIDEr EM
42.0 44.3 63.5 86.5 52.7
44.0 48.1 68.4 86.2 52.4
2.0 8.2 16.1 73.2 49.7
46.7 49.5 73.6 89.4 53.8
41.5 44.9 63.3 86.9 53.1
49.5 52.8 73.0 90.4 54.1
51.8 55.1 77.2 93.7 55.7
Table 2: Modality ablation results across five benchmarks. Text-only features underperform due to lack of geometric grounding, while 2D and 3D features alone offer moderate performance. The best results are achieved when all three modalities are integrated, highlighting their complementary strengths.
@@ 4.1 Datasets and Metrics
3DVG-Transformer [zhao20213dvg] - 25.5
3DJCG [cai20223djcg] - 26.6
D3Net [chen2022d3net] - 32.2
M3DRef-CLIP [zhang2023multi3drefer] 42.8 38.4
PQ3D [zhu2024unifying] - 50.1
@@
Grounded 3D-LLM [chen2024grounded] 44.7 40.8
Chat-Scene [huang2024chat] 57.1 52.4
@@
Table 3: Performance on Multi3DRefer [zhang2023multi3drefer]. Descrip3D achieves the best performance, with notable gains in multi-object reasoning due to its dual-level relational text modeling.
@@
Scan2Cap [chen2021scan2cap] 35.2 22.4
3DJCG [cai20223djcg] 49.5 31.0
3D-VLP [jin2023context] 54.9 32.3
3D-VisTA [zhu20233d] 66.9 34.0
Vote2Cap-DETR++ [chen2024vote2cap] 67.6 37.1
@@
LL3DA [chen2024ll3da] 65.2 36.8
LEO [huang2023embodied] 68.4 36.9
Grounded 3D-LLM [chen2024grounded] 70.2 35.0
Chat-Scene [huang2024chat] 77.1 36.3
@@
Table 4: Performance on Scan2Cap [chen2021scan2cap]. Descrip3D achieves the best C@0.5 performance by leveraging richer object-level relational context.
4.2 Implementation Details
We extract 100 object proposals per scene using the Mask3D [schult2022mask3d] segmentation model. For 3D geometry, we use Uni3D [zhou2023uni3d] to extract point embeddings $\mathbf{Z}^{v_{p}}_{i}$. For 2D appearance, we adopt DINOv2 [oquab2023dinov2] to extract object-level embeddings $\mathbf{Z}^{v_{f}}_{i}$, consistent with prior work (e.g., ChatScene [huang2024chat], 3DGraphLLM [zemskova20243dgraphllm]), to ensure fair comparison. While stronger encoders could raise performance, our relational-text modality is independent of the 2D backbone. For text, we utilize the LLaVA-v1.5-7B [liu2023llava] model to generate descriptions for each object. These object descriptions are then encoded using the all-mpnet-base-v2 model, a SentenceTransformer [reimers-2019-sentence-bert] based on MPNet [song2020mpnet] that captures rich contextual semantics for downstream fusion, resulting in the embeddings denoted as $\mathbf{Z}^{v_{t}}_{i}$. Each modality is projected to the token space using a three-layer MLP.
We use Vicuna-7B-v1.5 [chiang2023vicuna] as the base LLM to ensure a fair comparison with prior work (ChatScene, 3DGraphLLM, 3D-Ground-LLM). Grounded-3D-LLM primarily reports results with Tiny-Vicuna-1B, but also notes little difference with Vicuna-7B. We fine-tune it with LoRA [hu2022lora] (rank 16) for 3 epochs, batch size 32, lr $5\times 10^{-6}$, cosine schedule. Training is conducted on two 80G NVIDIA A100 GPUs and completes in about 24 hours. We observe that training for 2 epochs yields improved performance on the ScanQA dataset, and we adopt this setting in relevant evaluations.
4.3 Performance Comparison
@@
Scene-level Question Answering.
We evaluate our model on ScanQA [azuma2022scanqa] and SQA3D [ma2022sqa3d], both of which require comprehensive scene-level understanding and precise object-grounded reasoning. As shown in table˜1, our method achieves the highest scores across all major metrics compared with previous expert models and LLM-based methods, highlighting the strength of our dual-level integrated relational descriptions in enhancing context understanding. Notably, CIDEr, which emphasizes content relevance, shows a significant gain of 5.9 over PQ3D and 6.0 over Chat-Scene, confirming that our model generates more informative and relevant answers. On the SQA3D dataset, our method attains the highest EM and EM-R, again outperforming existing models. The strong results across both benchmarks demonstrate the effectiveness of enriching 3D object representations with object-level textual descriptions. As a general-purpose model, our model consistently outperforms task-specific systems, demonstrating strong versatility without the need for task-dependent architectures. Compared to knowledge-graph-based approaches such as 3DGraphLLM, our model achieves better alignment with language and context, suggesting that lightweight textual descriptions offer a more direct and interpretable semantic grounding. Overall, the results confirm that injecting fine-grained natural language descriptions of objects into the scene representation, through dual-level integration, significantly enhances the LLM’s ability to handle 3D visual question answering tasks.
3D Visual Grounding.
We evaluate 3D visual grounding on the ScanRefer [chen2020scanrefer] dataset. As shown in table˜1, our method achieves the best performance among all methods. Compared to strong expert models like ConcreteNet and 3D-VisTA, our model surpasses them by 5.3 and 6.3 at the stricter 0.5 IoU threshold, respectively. These expert models are trained with tailored 3D architectures and task-specific objectives, highlighting the strength of our unified and generalizable approach. Among LLM-based methods, our model also clearly outperforms Chat-Scene and 3DGraphLLM. The improvement over Chat-Scene, which uses only 2D and 3D embeddings, validates the great benefit of incorporating relational object-level textual descriptions through a dual-level integration approach.
Multi‑Object 3D Visual Grounding.
We evaluate our method on Multi3DRefer [zhang2023multi3drefer], a challenging benchmark requiring models to resolve complex multi-object visual grounding within 3D scenes. Unlike traditional single-object grounding, this task demands fine-grained relational reasoning among multiple entities. As shown in table˜3, our method achieves the best performance, outperforming both expert models and LLM-based baselines such as Chat-Scene and Grounded 3D-LLM.
3D Captioning.
On Scan2Cap [chen2021scan2cap], which evaluates the generation of 3D-aware captions for objects in complex indoor scenes, our model obtains the highest CIDEr score as shown in table˜4. This indicates that its captions are more informative and semantically aligned with ground truth. While BLEU-4 is slightly lower, this metric favors exact n-gram overlap and can penalize the diverse relational phrasing produced by our model. Importantly, unlike expert captioning systems tailored to this task, Descrip3D is trained as a unified model across multiple benchmarks yet still delivers strong captioning results. This shows that relational textual descriptions not only improve grounding and QA but also enhance semantic richness and generalization.
Qualitative results.
fig.˜3 presents qualitative comparisons between our method and Chat-Scene on two representative tasks: 3D question answering (fig.˜3(a)) and 3D visual grounding (fig.˜3(b)). For question answering, our model shows stronger ability to understand relational context and resolve spatial references, whereas Chat-Scene struggles more with object relationships. For visual grounding, our model more accurately identifies target objects based on dual-level integrated complex natural language descriptions involving appearance and relative position, while Chat-Scene often fails to disambiguate visually or semantically similar candidates. These results highlight the benefit of injecting detailed object-level descriptions into the language model through a dual-level approach, enabling more precise and interpretable multi-modal reasoning.
In summary, our method achieves strong and consistent results across five benchmarks, demonstrating its effectiveness across tasks.
4.4 Ablation Studies
To evaluate the impact of our object-level text descriptions and dual-level integration strategies, we perform ablations on key components of the model.
Effect of Individual Modalities.
We ablate 3D, 2D, and text features as shown in table˜2. Single modalities are limited: 3D captures geometry but lacks appearance and relations, 2D provides surface cues without spatial grounding, and text alone conveys semantics but cannot localize objects. Pairwise combinations yield complementary gains, while integrating all three produces the strongest results. This confirms that relational text is most effective when fused with both 2D and 3D features.
Effect of Object Reference Style in Prompt-Level Description.
We test different ways of expressing objects in prompt-level text injection as shown in table˜5. Using raw object names leads to ambiguity when multiple instances share the same category. Adding IDs to names introduces redundancy without improvement. In contrast, using IDs alone consistently gives the best results, ensuring unambiguous alignment between queries and object descriptions across all benchmarks.
Effect of Object Descriptions and Fusion Location.
We also study where textual descriptions are integrated, as shown in table˜6. Injecting them only as embeddings enriches object semantics and improves QA tasks. Injecting them only into prompts benefits tasks like Scan2Cap and Multi3DRefer. Combining both yields the best overall results across all benchmarks, with especially strong improvements on ScanQA. These findings show that embedding- and prompt-level strategies are complementary, and together provide the most robust improvements in both spatial grounding and semantic reasoning.
5 Conclusion